# The defensible set and a new impossibility theorem in voting

### 11 Jan 2024 v1

@@ Abstract
In the context of social choice theory with ordinal preferences, we say that the defensible set is the set of alternatives $x$ such that for any alternative $y$, if $y$ beats $x$ in a head-to-head majority comparison, then there is an alternative $z$ that beats $y$ in a head-to-head majority comparison by a margin at least as large as the margin by which $y$ beat $x$. We show that any ordinal voting method satisfying two well-known axioms from voting theory—positive involvement and the Condorcet winner criterion—refines the defensible set. Using this lemma, we prove an impossibility theorem: there is no such voting method that also satisfies the Condorcet loser criterion, resolvability, and a common invariance property for Condorcet methods, namely that the choice of winners depends only on the relative sizes of majority margins.
@@ 1 Introduction
In social choice based on ordinal preferences, there is an instability in choosing any alternative $x$ that loses a head-to-head majority comparison to another alternative $y$. The supporters of $y$ over $x$, being more numerous than those of $x$ over $y$, may claim on that basis that the choice of $x$ was illegitimate. They may advocate for the overthrow of $x$ in favor of $y$. However, we might defend the choice of $x$ on the following grounds, when applicable: there is another $z$ that beats $y$ head-to-head by a margin at least as large as that by which $y$ beat $x$; thus, the reason invoked to overthrow $x$ in favor of $y$ would immediately provide at least as strong a reason to overthrow $y$ in favor of $z$, undercutting the original idea to overthrow $x$ in favor of $y$. Indeed, there are always some alternatives $x$ that are defensible in this way.
For example, the alternatives selected by the Minimax (Simpson 1969, Kramer 1977) and Split Cycle (Holliday and Pacuit 2023a) voting methods are always so defensible. Minimax selects the alternatives whose worst head-to-head loss is smallest among any alternative’s worst head-to-head loss. Thus, if $x$ is a Minimax winner that loses head-to-head to $y$, there must be an alternative $z$ to which $y$ loses by at least as large a margin. As noted by Young (1977), the Minimax winner is the “most stable against overthrow” (p. 350). If $x$ is a Split Cycle winner that loses head-to-head to $y$, then there is a sequence of alternatives $y_{1},\dots,y_{n}$ with $y=y_{1}$ and $x=y_{n}$ such that each loses to the next by a margin at least as large as that by which $x$ lost to $y$; [^1] For the proof that there always exists such an $x$, see Holliday and Pacuit 2023a, Lemma 3.9. hence the reason for overthrowing $x$ in favor of $y$ would similarly justify a sequence of revolutions leading right back to $x$. By contrast, the Minimax winner $x$ might not be such that a sequence of revolutions as strong as the first would lead back to $x$. Nonetheless, the Minimax and Split Cycle winners are defensible against those who advocate an initial revolution. This kind of defensibility is equivalent to what Heitzig (2002) calls weak immunity to binary arguments (cf. Dung 1995) based on majority margins.
In this note, we relate defensibility to two well-known axioms in voting theory: positive involvement (Saari 1995) and the Condorcet winner criterion (Condorcet 1785). We show that any ordinal voting method satisfying these axioms chooses from among the defensible alternatives (Lemma 2). Using this lemma, we prove an impossibility theorem (Theorem 4): there is no such voting method that also satisfies the Condorcet loser criterion (Condorcet 1785), resolvability (Tideman 1986), and a common invariance property for Condorcet methods, namely that the choice of winners depends only on the relative sizes of majority margins. Methods satisfying this relative margin invariance include not only Minimax and Split Cycle but also Ranked Pairs (Tideman 1987), Beat Path (Schulze 2011), Weighted Covering (Dutta and Laslier 1999, Pérez-Fernández and De Baets 2018), and Stable Voting (Holliday and Pacuit 2023b). Thus, our theorem helps explain why none of these voting methods satisfies all the stated axioms, as well as motivating the search for new methods satisfying all the axioms except for relative margin invariance. [^2] A notebook with code to generate Table 1 and Figures 25 below is available at github.com/wesholliday/defensible-set.
2 Defensibility
@@
$Margin_{\mathbf{P}}(x,y)=|\{i\in V(\mathbf{P})\mid(x,y)\in\mathbf{P}(i)\}|-|\{i\in V(\mathbf{P})\mid(y,x)\in\mathbf{P}(i)\}|.$
A voting method is a function $F$ assigning to each profile $\mathbf{P}$ a nonempty $F(\mathbf{P})\subseteq X(\mathbf{P})$. [^5] Thus, we build the axiom of universal domain into the definition of our voting methods. If $|F(\mathbf{P})|>1$, we assume some further (possibly random) tiebreaking process ultimately narrows $F(\mathbf{P})$ down to one alternative.
@@
Given a profile $\mathbf{P}$, we define the defensible set of $\mathbf{P}$ as
$D(\mathbf{P})=\{x\in X(\mathbf{P})\mid\mbox{for all }y\in X\mbox{, there exists }z\in X:Margin_{\mathbf{P}}(z,y)\geq Margin_{\mathbf{P}}(y,x)\}.$
As a voting method, $D$ is a special case, based on majority margins, of Heitzig’s (2002) family of weakly immune methods. [^6] In particular, $D$ is the coarsest voting method satisfying all of Heitzig’s axioms $\mathrm{wIm}_{P_{\alpha}}$ for $\alpha\in(0,1]$. Equivalently, in the argumentation theory of Dung (1995, Definition 6.1), the alternatives in $D(\mathbf{P})$ are acceptable with respect to $X(\mathbf{P})$ given any attack relation $attacks_{n}$ for any positive integer $n$ defined by $attacks_{n}(x,y)$ if $Margin_{\mathbf{P}}(x,y)\geq n$. In practice, we are interested in methods that are more resolute than $D$, in the sense that they pick multiple alternatives less frequently (a point discussed below). A voting method $F$ refines the defensible set (for some class of profiles) if for any profile $\mathbf{P}$ (in the class), $F(\mathbf{P})\subseteq D(\mathbf{P})$. Presently we will show that any method $F$ satisfying two appealing axioms refines the defensible set.
$F$ satisfies positive involvement (see Saari 1995, Pérez 2001, Holliday and Pacuit 2021) if for any profiles $\mathbf{P},\mathbf{P}^{\prime}$, if $x\in F(\mathbf{P})$ and $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding one voter who ranks $x$ uniquely in first place, then $x\in F(\mathbf{P}^{\prime})$. [^7] The requirement that the new voter ranks $x$ uniquely first is crucial; a stronger version of the axiom that applies whenever the new voter does not strictly prefer any $y$ to $x$ is inconsistent with the Condorcet winner criterion (Pérez 2001, Duddy 2014). Thus, you cannot cause $x$ to lose by ranking $x$ first. According to Pérez (2001, p. 605), this “may be seen as the minimum to require concerning the coherence in the winning set when new voters are added.”
$F$ satisfies the Condorcet winner criterion (Condorcet 1785) if for every profile $\mathbf{P}$ with a Condorcet winner, $F(\mathbf{P})$ contains only the Condorcet winner. Recall that $x$ is a Condorcet winner (resp. weak Condorcet winner) if $x$ beats (resp. does not lose to) any other alternative head-to-head, i.e., for every $y\in X(\mathbf{P})\setminus\{x\}$, $Margin_{\mathbf{P}}(x,y)>0$ (resp. $Margin_{\mathbf{P}}(x,y)\geq 0$). $F$ satisfies the weak Condorcet winner criterion if for every profile $\mathbf{P}$ with a weak Condorcet winner, $F(\mathbf{P})$ contains only weak Condorcet winners.
@@
For part 1, let $F$ be such a method, $\mathbf{P}$ a linear profile, and $x\in F(\mathbf{P})$. Toward a contradiction, suppose $x\not\in D(\mathbf{P})$. Hence there is a $y\in X$ with $Margin_{\mathbf{P}}(y,x)>0$ and for every $z\in X$, $Margin_{\mathbf{P}}(z,y)<Margin_{\mathbf{P}}(y,x)$. Let $k=max\{Margin_{\mathbf{P}}(z,y)\mid z\in X(\mathbf{P})\}$, so $k<Margin_{\mathbf{P}}(y,x)$. Since $\mathbf{P}$ is linear, all margins have the same parity, so $k+1<Margin_{\mathbf{P}}(y,x)$. Let $\mathbf{P}^{\prime}$ be obtained from $\mathbf{P}$ by adding $k+1$ voters who rank $x$ uniquely first and $y$ uniquely second, followed by any linear order of the remaining alternatives. Then by positive involvement, $x\in F(\mathbf{P}^{\prime})$. But in $\mathbf{P}^{\prime}$, $y$ is the Condorcet winner, so by the Condorcet winner criterion, $F(\mathbf{P}^{\prime})=\{y\}$, contradicting $x\in F(\mathbf{P}^{\prime})$. Thus, $x\in D(\mathbf{P})$.
For part 2 and an arbitrary $\mathbf{P}$, the argument is similar only the margins may have different parities, so we cannot infer from $k<Margin_{\mathbf{P}}(y,x)$ that $k+1<Margin_{\mathbf{P}}(y,x)$. In this case, we add only $k$ voters who rank $x$ uniquely first and $y$ uniquely second, followed by any linear order of the remaining alternatives. Then in $\mathbf{P}^{\prime}$, $y$ is a weak Condorcet winner, and $Margin_{\mathbf{P}^{\prime}}(y,x)>0$. It follows by the weak Condorcet winner criterion that $x\not\in F(\mathbf{P}^{\prime})$, contradicting positive involvement.∎
Given the appeal of positive involvement and the Condorcet winner criterion, it is of significant interest to explore refinements of the defensible set. Minimax and Split Cycle satisfy both axioms (Pérez 2001, Holliday and Pacuit 2023a), which again shows they refine the defensible set. However, refinements of the defensible set may violate positive involvement. Examples include some refinements of Split Cycle such as Beat Path, Ranked Pairs, and Stable Voting (see Holliday and Pacuit 2023a).
@@ 3 Impossibility
Let us now try adding further axioms to our wish list. A Condorcet loser in a profile $\mathbf{P}$ is an $x\in X(\mathbf{P})$ that loses head-to-head to every other alternative, i.e., for every $y\in X(\mathbf{P})\setminus\{x\}$, $Margin_{\mathbf{P}}(y,x)>0$. $F$ satisfies the Condorcet loser criterion if for every profile $\mathbf{P}$, $F(\mathbf{P})$ does not contain a Condorcet loser.
@@
While the normative appeal of positive involvement and the Condorcet criteria, as well as the practical relevance of resolvability, is clear, it is less obvious whether relative margin invariance should be a desideratum. One point in favor of methods satisfying relative margin invariance is that a small amount of noise in the collection of voter preferences is unlikely to change the relative margin graph and therefore unlikely to change the selection of winners, rendering such rules quite robust to noise (cf. Procaccia et al. 2006).
@@
If $F$ satisfies relative margin invariance and single-voter (resp. asymptotic) resolvability, then $F$ selects a unique winner in any profile whose relative margin graph is a linearly edged-ordered tournament.
@@
For single-voter resolvability, suppose $F$ selects multiple winners in a profile $\mathbf{P}$ whose relative margin graph is a linearly edge-ordered tournament. Consider $4\mathbf{P}$, the profile obtained from $\mathbf{P}$ by replacing each voter with $4$ copies of that voter. Since $4\mathbf{P}$ has the same relative margin graph as $\mathbf{P}$, $F$ selects the same winners in $4\mathbf{P}$ as in $\mathbf{P}$. But by adding a single voter to $4\mathbf{P}$, it is impossible to obtain a $\mathbf{P}^{\prime}$ whose relative margin graph differs from that of $4\mathbf{P}$, so it is impossible to obtain a profile with a unique winner. Hence $F$ does not satisfy single-voter resolvability.
For asymptotic resolvability, Harrison-Trainor (2022, Theorem 11.2) shows that for every positive integer $m$ and linearly edge-ordered tournament $T$ for $m$ alternatives, in the limit as the number of voters goes to infinity, the proportion of linear profiles for $m$ alternatives whose relative margin graph is $T$ is nonzero. Thus, if $F$ picks multiple winners in such a tournament, $F$ does not satisfy asymptotic resolvability.∎
Thanks to Lemma 3, either version of resolvability works below, so we simply speak of ‘resolvability’. The defensible set does not satisfy either version, since it returns multiple winners for some linearly edge-ordered tournaments. Table 1 shows the extent of its irresoluteness for linearly edge-ordered tournaments for four alternatives in comparison to several other voting methods. [^8] The Smith set (Smith 1973) is the smallest nonempty set of alternatives such that every alternative in the set beats ever alternative outside the set head-to-head. The uncovered set (Fishburn 1977, Miller 1980) is the set of all alternatives $x$ for which there is no $y$ such that $y$ beats $x$ and beats every $z$ that $x$ beats; this definition is equivalent to others (see Duggan 2013) provided there are no zero margins between distinct alternatives, which is the case for Table 1. In this case, the Copeland winners (Copeland 1951) are the alternatives that beat the most other alternatives.
@@
Split Cycle $104$ $1.0541\overline{6}$ $2$
Minimax 0 1 $1$
@@
If we drop any of the axioms introduced so far besides the Condorcet winner criterion and relative margin invariance, there is a method satisfying the remaining axioms: dropping positive involvement, Ranked Pairs works; dropping the Condorcet loser criterion, Minimax works; [^9] Indeed, it is plausible that with more combinatorial work, our proof strategy for Theorem 4 could yield a characterization of Minimax for four alternatives using the axioms other than the Condorcet loser criterion (cf. Holliday and Pacuit 2023c). and dropping resolvability, Split Cycle works. If we drop the Condorcet winner criterion and relative margin invariance, then the Borda method (Borda 1781) works. However, no method satisfies all of the axioms.
@@
There is no voting method satisfying positive involvement, the Condorcet winner criterion, the Condorcet loser criterion, resolvability, and relative margin invariance.
@@
Assume there is such an $F$. To derive a contradiction, we use the relative margin graphs in Figure 1. The numbers indicate the ordering $\succ$ from the smallest margin ($1$) to the largest ($6$). The defensible set for each graph is shaded in gray. Since $F$ satisfies relative margin invariance and resolvability, $F$ returns a singleton for each graph by Lemma 3. Given $\mathbb{M}$ and $\mathbb{M}^{\prime}$ and an alternative $x$, we write $\mathbb{M}\Rightarrow_{x}\mathbb{M}^{\prime}$ in Figure 1 if there are profiles $\mathbf{P},\mathbf{P}^{\prime}$ such that $\mathbb{M}$ is the relative margin graph of $\mathbf{P}$, $\mathbb{M}^{\prime}$ is the relative margin graph of $\mathbf{P}^{\prime}$, and $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding voters all of whom rank $x$ uniquely first. The construction of such profiles is an integer linear programming problem, whose solution yields the profiles in Figures 25.
@@
Figure 1: Relative margin graphs for the proof of Theorem 4.
Suppose $F(\mathbb{M}_{1})=\{d\}$. Then by positive involvement and resolvability, $F(\mathbb{M}_{2})=\{d\}$. On the other hand, $F(\mathbb{M}_{3})\subseteq\{b,d\}$ by Lemma 2, but $d\not\in F(\mathbb{M}_{3})$ by the Condorcet loser criterion, so $F(\mathbb{M}_{3})=\{b\}$. Then by positive involvement and resolvability, $F(\mathbb{M}_{2})=\{b\}$, contradicting $F(\mathbb{M}_{2})=\{d\}$.
Thus, $F(\mathbb{M}_{1})=\{a\}$. Then by positive involvement and resolvability, $F(\mathbb{M}_{4})=\{a\}$. But $F(\mathbb{M}_{5})=\{d\}$ by Lemma 2, so by positive involvement and resolvability, $F(\mathbb{M}_{4})=\{d\}$, contradicting $F(\mathbb{M}_{4})=\{a\}$.∎
Thus, in the search for the “holy grail” of a voting method satisfying positive involvement, the Condorcet winner and loser criteria, and resolvability, we must drop the restriction of relative margin invariance. Whether such a voting method exists or another impossibility theorem awaits us is an important open question. We believe that the defensible set will likely be useful in finding the answer.

### 26 Feb 2024 v4

# An impossibility theorem concerning positive involvement in voting

@@ Abstract
In social choice theory with ordinal preferences, a voting method satisfies the axiom of positive involvement if adding to a preference profile a voter who ranks an alternative uniquely first cannot cause that alternative to go from winning to losing. In this note, we prove a new impossibility theorem concerning this axiom: there is no ordinal voting method satisfying positive involvement that also satisfies the Condorcet winner and loser criteria, resolvability, and a common invariance property for Condorcet methods, namely that the choice of winners depends only on the ordering of majority margins by size.
@@ 1 Introduction
A basic assumption of democratic voting is that increased voter support for a candidate should not harm that candidate’s chances of winning. In social choice theory with ordinal preferences, one formalization of this idea is given by the axiom of positive involvement (Saari 1995): adding to a preference profile a voter who ranks an alternative uniquely first cannot cause that alternative to go from winning to losing. [^1] Of course, such a voter expresses more than just support for their favorite alternative if they also rank other alternatives. A weaker formalization is therefore bullet-vote positive involvement: adding a voter who ranks an alternative uniquely first and all other alternatives in a tie below their favorite cannot cause their favorite to go from winning to losing. Perhaps surprisingly, many well-known voting methods—especially Condorcet-consistent voting methods—violate this axiom (see Pérez 1995). In this note, we prove a new impossibility theorem in this vein: there is no ordinal voting method satisfying positive involvement that also satisfies the Condorcet winner and loser criteria (Condorcet 1785), resolvability (Tideman 1986), and a common invariance property for Condorcet methods, namely that the choice of winners depends only on the ordering of majority margins by size. Methods satisfying this ordinal margin invariance include Minimax (Simpson 1969, Kramer 1977), Ranked Pairs (Tideman 1987), Beat Path (Schulze 2011), Weighted Covering (Dutta and Laslier 1999, Pérez-Fernández and De Baets 2018), Split Cycle (Holliday and Pacuit 2023a), and Stable Voting (Holliday and Pacuit 2023b). Thus, our theorem helps explain why none of these voting methods satisfies all the stated axioms, as well as motivating the search for new methods satisfying all the axioms except for ordinal margin invariance.
2 Preliminaries
@@
$\text{Margin}_{\mathbf{P}}(x,y)=\text{Support}_{\mathbf{P}}(x,y)-\text{Support}_{\mathbf{P}}(y,x)\mbox{, where }\text{Support}_{\mathbf{P}}(a,b)=|\{i\in V(\mathbf{P})\mid(a,b)\in\mathbf{P}(i)\}|.$
A voting method is a function $F$ assigning to each profile $\mathbf{P}$ a nonempty $F(\mathbf{P})\subseteq X(\mathbf{P})$. [^4] Thus, we build the axiom of universal domain into the definition of our voting methods. If $|F(\mathbf{P})|>1$, we assume some further (possibly random) tiebreaking process ultimately narrows $F(\mathbf{P})$ down to one alternative. A voting method $G$ refines $F$ (for some class of profiles) if for any profile $\mathbf{P}$ (in the class), $G(\mathbf{P})\subseteq F(\mathbf{P})$.
A voting method $F$ satisfies positive involvement (see Saari 1995, Pérez 2001, Holliday and Pacuit 2021) if for any profiles $\mathbf{P},\mathbf{P}^{\prime}$, if $x\in F(\mathbf{P})$ and $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding one voter who ranks $x$ uniquely in first place, then $x\in F(\mathbf{P}^{\prime})$. [^5] The requirement that the new voter ranks $x$ uniquely first is crucial; a stronger version of the axiom that applies whenever the new voter does not strictly prefer any $y$ to $x$ is inconsistent with the Condorcet winner criterion (Pérez 2001, Duddy 2014). Thus, you cannot cause $x$ to lose by ranking $x$ first. According to Pérez (2001, p. 605), this “may be seen as the minimum to require concerning the coherence in the winning set when new voters are added.”
A voting method $F$ satisfies the Condorcet winner criterion (Condorcet 1785) if for every profile $\mathbf{P}$ with a Condorcet winner, $F(\mathbf{P})$ contains only the Condorcet winner. Recall that $x$ is a Condorcet winner (resp. weak Condorcet winner) if $x$ beats (resp. does not lose to) any other alternative head-to-head, i.e., for every $y\in X(\mathbf{P})\setminus\{x\}$, $\text{Margin}_{\mathbf{P}}(x,y)>0$ (resp. $\text{Margin}_{\mathbf{P}}(x,y)\geq 0$). $F$ satisfies the weak Condorcet winner criterion if for every profile $\mathbf{P}$ with a weak Condorcet winner, $F(\mathbf{P})$ contains only weak Condorcet winners.
A key tool for reasoning about voting methods satisfying both positive involvement and the Condorcet winner criterion is given by the following voting method, appearing implicitly in Lemma 3 of Pérez 1995 and studied explicitly in Kasper et al. 2019.
@@
Given a profile $\mathbf{P}$, define the defensible set of $\mathbf{P}$ as
$D(\mathbf{P})=\{x\in X(\mathbf{P})\mid\mbox{for all }y\in X(\mathbf{P})\mbox{,there exists }z\in X(\mathbf{P}):\text{Margin}_{\mathbf{P}}(z,y)\geq\text{Margin}_{\mathbf{P}}(y,x)\}.$
As a voting method, $D$ is a special case, based on majority margins, of Heitzig’s (2002) family of methods that are weakly immune to binary arguments. [^6] In particular, $D$ is the coarsest voting method satisfying all of Heitzig’s axioms $\mathrm{wIm}_{P_{\alpha}}$ for $\alpha\in(0,1]$. The idea can also be found in the argumentation theory of Dung (1995). [^7] In Dung’s (1995, Definition 6.1) terms, the alternatives in $D(\mathbf{P})$ are acceptable with respect to $X(\mathbf{P})$ given each attack relation $attacks_{n}$, for each positive integer $n$, defined by $attacks_{n}(x,y)$ if $\text{Margin}_{\mathbf{P}}(x,y)\geq n$. Our choice of the term ‘defensible’ is based on the following intuition. Although $x$ may lose head-to-head to $y$, possibly prompting the supporters of $y$ over $x$ to call for the overthrow of $x$ in favor of $y$, we can defend the choice of $x$ on the following grounds, when applicable: there is another $z$ that beats $y$ head-to-head by a margin at least as large as that by which $y$ beat $x$; thus, the reason invoked to overthrow $x$ in favor of $y$ would immediately provide at least as strong a reason to overthrow $y$ in favor of $z$, undercutting the original idea to overthrow $x$ in favor of $y$.
Part 1 of the following is equivalent to Lemma 3 of Pérez 1995 (cf. Lemma 1 of Pérez 2001), [^8] For linear profiles, where $\text{Support}_{\mathbf{P}}(x,y)+\text{Support}_{\mathbf{P}}(y,x)=|V(\mathbf{P})|$, one can show that $\text{Margin}_{\mathbf{P}}(z,y)\geq\text{Margin}_{\mathbf{P}}(y,x)$ is equivalent to $\text{Support}(x,y)\geq\text{Support}(y,z)$, which is the relevant condition in Lemma 3 of Pérez 1995. which is in turn an adaptation of a similar result of Moulin 1988, but we include a proof to keep this note self-contained.
@@
For part 1, let $F$ be such a method, $\mathbf{P}$ a linear profile, and $x\in F(\mathbf{P})$. Toward a contradiction, suppose $x\not\in D(\mathbf{P})$. Hence there is a $y\in X(\mathbf{P})$ with $\text{Margin}_{\mathbf{P}}(y,x)>0$ and for every $z\in X(\mathbf{P})$, $\text{Margin}_{\mathbf{P}}(z,y)<\text{Margin}_{\mathbf{P}}(y,x)$. Let $k=max\{\text{Margin}_{\mathbf{P}}(z,y)\mid z\in X(\mathbf{P})\}$, so $k<\text{Margin}_{\mathbf{P}}(y,x)$. Since $\mathbf{P}$ is linear, all margins have the same parity, so $k+1<\text{Margin}_{\mathbf{P}}(y,x)$. Let $\mathbf{P}^{\prime}$ be obtained from $\mathbf{P}$ by adding $k+1$ voters who rank $x$ uniquely first and $y$ uniquely second, followed by any linear order of the remaining alternatives. Then by positive involvement, $x\in F(\mathbf{P}^{\prime})$. But in $\mathbf{P}^{\prime}$, $y$ is the Condorcet winner, so by the Condorcet winner criterion, $F(\mathbf{P}^{\prime})=\{y\}$, contradicting $x\in F(\mathbf{P}^{\prime})$. Thus, $x\in D(\mathbf{P})$.
For part 2 and an arbitrary $\mathbf{P}$, the argument is similar only the margins may have different parities, so we cannot infer from $k<\text{Margin}_{\mathbf{P}}(y,x)$ that $k+1<\text{Margin}_{\mathbf{P}}(y,x)$. In this case, we add only $k$ voters who rank $x$ uniquely first and $y$ uniquely second, followed by any linear order of the remaining alternatives. Then in $\mathbf{P}^{\prime}$, $y$ is a weak Condorcet winner, and $\text{Margin}_{\mathbf{P}^{\prime}}(y,x)>0$. It follows by the weak Condorcet winner criterion that $x\not\in F(\mathbf{P}^{\prime})$, contradicting positive involvement.∎
Given the appeal of positive involvement and the Condorcet winner criterion, it is of significant interest to explore refinements of the defensible set (cf. Kasper et al. 2019, § 4). The Minimax and Split Cycle methods satisfy these axioms (Pérez 2001, Holliday and Pacuit 2023a), so they refine the defensible set. [^9] This is also clear directly. Minimax selects the alternatives whose worst head-to-head loss is smallest among any alternative’s worst head-to-head loss. Thus, if $x$ is a Minimax winner that loses head-to-head to $y$, there must be an alternative $z$ to which $y$ loses by at least as large a margin. As noted by Young (1977), the Minimax winner is the “most stable against overthrow” (p. 350). If $x$ is a Split Cycle winner that loses head-to-head to $y$, then there is a sequence of alternatives $y_{1},\dots,y_{n}$ with $y=y_{1}$ and $x=y_{n}$ such that each loses to the next by a margin at least as large as that by which $x$ lost to $y$; hence the reason for overthrowing $x$ in favor of $y$ would similarly justify a sequence of revolutions leading right back to $x$. However, refinements of the defensible set may violate positive involvement. Examples include some refinements of Split Cycle such as Beat Path, Ranked Pairs, and Stable Voting (see Holliday and Pacuit 2023a).
@@ 3 Impossibility
Let us now try adding further axioms to our wish list. A Condorcet loser in a profile $\mathbf{P}$ is an $x\in X(\mathbf{P})$ that loses head-to-head to every other alternative, i.e., for every $y\in X(\mathbf{P})\setminus\{x\}$, $\text{Margin}_{\mathbf{P}}(y,x)>0$. $F$ satisfies the Condorcet loser criterion if for every profile $\mathbf{P}$, $F(\mathbf{P})$ does not contain a Condorcet loser.
@@
While the normative appeal of positive involvement and the Condorcet criteria, as well as the practical relevance of resolvability, is clear, it is less obvious whether ordinal margin invariance should be a desideratum. One point in favor of methods satisfying ordinal margin invariance is that a small amount of noise in the collection of voter preferences is unlikely to change the ordinal margin graph and therefore unlikely to change the selection of winners, rendering such rules quite robust to noise (cf. Procaccia et al. 2006).
@@
If $F$ satisfies ordinal margin invariance and single-voter (resp. asymptotic) resolvability, then $F$ selects a unique winner in any profile whose ordinal margin graph is a linearly edged-ordered tournament.
@@
For single-voter resolvability, suppose $F$ selects multiple winners in a profile $\mathbf{P}$ whose ordinal margin graph is a linearly edge-ordered tournament. Consider $3\mathbf{P}$, the profile obtained from $\mathbf{P}$ by replacing each voter with three copies of that voter. Since $3\mathbf{P}$ has the same ordinal margin graph as $\mathbf{P}$, $F$ selects the same winners in $3\mathbf{P}$ as in $\mathbf{P}$. But by adding a single voter to $3\mathbf{P}$, it is impossible to obtain a $\mathbf{P}^{\prime}$ whose ordinal margin graph differs from that of $3\mathbf{P}$, so it is impossible to obtain a profile with a unique winner. Hence $F$ does not satisfy single-voter resolvability.
For asymptotic resolvability, Harrison-Trainor (2022, Theorem 11.2) shows that for every positive integer $m$ and linearly edge-ordered tournament $T$ for $m$ alternatives, in the limit as the number of voters goes to infinity, the proportion of linear profiles for $m$ alternatives whose ordinal margin graph is $T$ is nonzero. Thus, if $F$ picks multiple winners in such a tournament, $F$ does not satisfy asymptotic resolvability.∎
Thanks to Lemma 3, either version of resolvability works below, so we simply speak of ‘resolvability’. The defensible set does not satisfy either version, since it returns multiple winners for some linearly edge-ordered tournaments. Table 1 shows the extent of its irresoluteness for linearly edge-ordered tournaments for four alternatives in comparison to several other voting methods. [^10] The Smith set (Smith 1973) is the smallest nonempty set of alternatives such that every alternative in the set beats ever alternative outside the set head-to-head. The uncovered set (Fishburn 1977, Miller 1980) is the set of all alternatives $x$ for which there is no $y$ such that $y$ beats $x$ and beats every $z$ that $x$ beats; this definition is equivalent to others (see Duggan 2013) provided there are no zero margins between distinct alternatives, which is the case for Table 1. In this case, the Copeland winners (Copeland 1951) are the alternatives that beat the most other alternatives. A notebook with code to generate Table 1 and Figures 25 below is available at github.com/wesholliday/pos-inv.
@@
Split Cycle 104 $1.0541\overline{6}\phantom{1}$ 2
Minimax 110 1 1
@@
If we drop any of the axioms introduced so far besides the Condorcet winner criterion and ordinal margin invariance, then there is a method satisfying the remaining axioms. This is shown by Beat Path/Ranked Pairs/Stable Voting, Minimax, [^11] It is plausible that with more combinatorial work, our proof strategy for Theorem 4 could yield a characterization of Minimax for four alternatives using the axioms other than the Condorcet loser criterion (cf. Holliday and Pacuit 2023c). and Split Cycle in Table 2, which also includes the Borda method (Borda 1781) and Black’s method [^12] Black’s method selects the Condorcet winner only, if one exists, and otherwise selects all Borda winners. (Black 1958) for comparison. However, no method satisfies all of the axioms. [^13] In Theorem 4, positive involvement can be replaced by negative involvement (see Pérez 2001) by the proof of Proposition 3.19 in Ding et al. 2023, since ordinal margin invariance implies the neutral reversal axiom in that proposition.
Beat Path Black’s Borda Minimax Split Cycle
Ranked Pairs
Stable Voting
Positive involvement -- -- $\checkmark$ $\checkmark$ $\checkmark$
Condorcet winner $\checkmark$ $\checkmark$ -- $\checkmark$ $\checkmark$
Condorcet loser $\checkmark$ $\checkmark$ $\checkmark$ -- $\checkmark$
Resolvability $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ --
Ordinal margin invariance $\checkmark$ -- -- $\checkmark$ $\checkmark$
Table 2: Satisfaction ($\checkmark$) or violation (--) of the axioms by selected voting methods.
@@
There is no voting method satisfying positive involvement, the Condorcet winner criterion, the Condorcet loser criterion, resolvability, and ordinal margin invariance.
@@
Assume there is such an $F$. To derive a contradiction, we use the ordinal margin graphs in Figure 1. The numbers indicate the ordering $\succ$ from the smallest margin ($1$) to the largest ($6$). The defensible set for each graph is shaded in gray. Since $F$ satisfies ordinal margin invariance and resolvability, $F$ returns a singleton for each graph by Lemma 3. Given $\mathbb{M}$ and $\mathbb{M}^{\prime}$ and an alternative $x$, we write $\mathbb{M}\Rightarrow_{x}\mathbb{M}^{\prime}$ in Figure 1 if there are profiles $\mathbf{P},\mathbf{P}^{\prime}$ such that $\mathbb{M}$ is the ordinal margin graph of $\mathbf{P}$, $\mathbb{M}^{\prime}$ is the ordinal margin graph of $\mathbf{P}^{\prime}$, and $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding voters all of whom rank $x$ uniquely first. The construction of such profiles is an integer linear programming problem, whose solution yields the profiles in Figures 25.141414Although we minimize the number of voters in these profiles subject to the relevant constraints, this does not answer an interesting question: what is the minimal number of voters needed for the impossibility theorem itself (cf. Brandt et al. 2017)?
Suppose $F(\mathbb{M}_{1})=\{d\}$. Then by positive involvement and resolvability, $F(\mathbb{M}_{2})=\{d\}$. On the other hand, $F(\mathbb{M}_{3})\subseteq\{b,d\}$ by Lemma 2, but $d\not\in F(\mathbb{M}_{3})$ by the Condorcet loser criterion, so $F(\mathbb{M}_{3})=\{b\}$. Then by positive involvement, $b\in F(\mathbb{M}_{2})$, contradicting $F(\mathbb{M}_{2})=\{d\}$.
Thus, $F(\mathbb{M}_{1})=\{a\}$. Then by positive involvement and resolvability, $F(\mathbb{M}_{4})=\{a\}$. But $F(\mathbb{M}_{5})=\{d\}$ by Lemma 2, so by positive involvement, $d\in F(\mathbb{M}_{4})$, contradicting $F(\mathbb{M}_{4})=\{a\}$.∎
Thus, in the search for the “holy grail” of a voting method satisfying positive involvement, the Condorcet winner and loser criteria, and resolvability, we must drop the restriction of ordinal margin invariance. [^15] Natural weakenings of ordinal margin invariance to consider are the C1.5 (De Donder et al. 2000) and C2 (Fishburn 1977) invariance conditions. Whether such a method exists or another impossibility theorem awaits us is an important open question.
@@
Figure 1: Ordinal margin graphs for the proof of Theorem 4.